Factor
5xSQUARED-13x+6
could u show me how u got this please

Answer by stanbon(75887)

5xSQUARED-13x+6
-------
Think of two number whose product is 5*6=30
and whose sum is -13.
The numbers are -3 and -10
Replace -13x with the terms -3x and -10x to get:
5x^2-3x-10x+6=0
Factor the 1st two and the last two terms separately to get:
x(5x-3)-2(5x-3)=0
(5x-3)(x-2)=0
x=3/5 or x=2
============
Cheers,
Stan H.